# TRACT: Regression-Aware Fine-tuning Meets Chain-of-Thought Reasoning for LLM-as-a-Judge

### v1

@@ Abstract
The LLM-as-a-judge paradigm uses large language models (LLMs) for automated text evaluation, where a numerical assessment is assigned by an LLM to the input text following scoring rubrics. Existing methods for LLM-as-a-judge use cross-entropy (CE) loss for fine-tuning, which neglects the numeric nature of score prediction. Recent work addresses numerical prediction limitations of LLM fine-tuning through regression-aware fine-tuning, which, however, does not consider chain-of-thought (CoT) reasoning for score prediction. In this paper, we introduce TRACT (Two-stage Regression-Aware fine-tuning with CoT), a method combining CoT reasoning with regression-aware training. TRACT consists of two stages: first, seed LLM is fine-tuned to generate CoTs, which serve as supervision for the second stage fine-tuning. The training objective of TRACT combines the CE loss for learning the CoT reasoning capabilities, and the regression-aware loss for the score prediction. Experiments across four LLM-as-a-judge datasets and two LLMs show that TRACT significantly outperforms existing methods. Extensive ablation studies validate the importance of each component in TRACT. [^1] All code and models can be found at https://github.com/d223302/TRACT.
@@
Cheng-Han Chiang National Taiwan University Taiwan dcml0714@gmail.com                        Hung-yi Lee National Taiwan University Taiwan hungyilee@ntu.edu.tw                        Michal Lukasik Google Research New York mlukasik@google.com
@@ 1 Introduction
Large Language Models (LLMs) have been applied to evaluate written texts based on the evaluation criteria specified in the input, outputting a score indicating the assessed quality (Chiang and Lee, 2023a). This LLM-as-a-judge paradigm has become the standard way to evaluate LLMs (Zheng et al., 2023; Dubois et al., 2024; Abdin et al., 2024; Lambert et al., 2025), and has applications across diverse scenarios (Yuan et al., 2024; Chiang et al., 2024). To induce the fine-grained assessment ability of LLMs, LLM-as-a-judge is typically trained with cross-entropy (CE) loss to predict chain-of-thought (CoT) reasoning about the evaluation followed by a score (Kim et al., 2024a, b; Li et al., 2024).
@@
Although fine-tuning LLM as a judge with CE loss is a well-established practice (Kim et al., 2024a, b), it has intuitive limitations when applied to predicting numerical targets. For example, given a ground truth score ’1’, placing high probability on the token ’5’ is penalized the same as placing high probability on the token ’2’, despite the fact that they induce very different errors in terms of numerical metrics such as squared error. To mitigate this, Lukasik et al. (2025) recently proposed regression-aware fine-tuning (RAFT; see §2.3), a method using squared error loss in fine-tuning LLMs.
RAFT has shown promising results on regression tasks; however, it does not incorporate CoT reasoning Wei et al. (2022); Kojima et al. (2022). Chiang and Lee (2023b) show that not using CoT can harm the performance of LLM-as-a-judge. This prompts us to ask: Can CoT reasoning be effectively integrated into regression-aware fine-tuning?
In this paper, we find a positive answer to the above research question, and as a result introduce Two-stage Regression-Aware fine-tuning with CoT reasoning (TRACT), a method that combines the benefits of CoT reasoning with RAFT for improved numerical prediction abilities. TRACT’s fine-tuning objective (which we refer to as CoT-RAFT) is the sum of CE loss over the CoT supervision, and the RAFT loss over the ground truth scores (see Figure 1(a)). The key aspect of TRACT for achieving strong results is aligning the CoT supervision with the model’s CoT distribution, thus leading to two-stage fine-tuning algorithm, where both stages rely on the CoT-RAFT fine-tuning objective (see Figure 1(b)). In stage 1, TRACT fine-tunes seed LLM over both the CoT supervision and the ground truth scores in the original training data. In stage 2, TRACT fine-tunes seed LLM from the CoTs generated by the model trained in stage 1. We summarize TRACT in Algorithm 1 and contrast its formulation against baselines in Table 1.
We conduct experiments on four datasets and two LLMs and show that TRACT consistently outperforms baselines by a large margin, including Prometheus-2-7B (Kim et al., 2024b), which is the best model of equal size from prior works. Our contributions are as follows:
@@ 3.2 Regression-Aware Fine-Tuning with CoT
CoT reasoning elicits a reasoning $s$ before outputting the score $y$. We propose and define the CR (CoT-RAIL) predictor as first sampling a CoT $\hat{s}$ conditioning on the input $x$, and then applying the RAIL predictor when conditioning on both $x$ and $\hat{s}$,
@@ 3.3 Self-Generated Chain-of-Thoughts (CoTs)
When training with CoT-RAFT (Equation 4), the score predictor is conditioned on the CoT generated by a target model, denoted as $\hat{s}\sim p_{\rm t}(\cdot\,|\,x)$. However, at inference time, CoT-RAIL (Equation 3) relies on self-generated CoTs from the fine-tuned model $p$, i.e., $\hat{s}\sim p(\cdot\,|\,x)$. In our case, the CoTs used for training are from GPT-4 ($p_{\rm a}$) and may be very different from the CoTs generated by the fine-tuned model $p$. This creates a mismatch between the CoTs used during training and inference, which can harm the performance, as empirically shown in §5.
@@ 4 Experiment Setup
Models
We evaluate TRACT using two LLMs, Mistral-7B-Instruct-v0.2 (Jiang et al., 2023) and Llama-3.1-8B-Instruct (Grattafiori et al., 2024).
Training Dataset
We follow Kim et al. (2024b) and use Feedback Collection (Kim et al., 2024a) as the training set. The training set contains roughly 100K samples. The whole dataset, including the samples to be evaluated, the evaluation responses, and ground truth scores, are generated by GPT-4. We follow the procedure in §3.3 to construct $D_{\rm self}$. We analyze the quality of the self-generated CoTs in Appendix B.1.
Test Datasets
We use four test datasets, following Kim et al. (2024b).
@@
Baselines
We compare TRACT against multiple baselines that differ in their training objectives and in whether they use CoT reasoning, including: (1) Standard fine-tuning and decoding without CoT2.2), (2) Standard fine-tuning and decoding with CoT3.1), (3) Zero-shot RAIL: Using RAIL predictor (Equation 1) on the seed LLM $p_{0}$, (4) RAFT2.3). A detailed comparison of the baselines is shown in Table 1.
@@
Id CoT Train Data Inf. $r$ $\rho$ $r$ $\rho$ $r$ $\rho$ $r$ $\rho$ $r$ $\rho$
Baselines
B.1 CE GPT-4 Decode 0.890 0.891 0.355 0.361 0.429 0.414 0.279 0.268 0.488 0.483
B.2 CE GPT-4 Decode 0.872 0.872 0.413 0.407 0.463 0.456 0.480 0.482 0.557 0.554
B.3 RAIL 0.197 0.175 0.200 0.149 0.281 0.165 0.309 0.216 0.247 0.176
B.4 RAFT GPT-4 RAIL 0.932 0.930 0.509 0.502 0.567 0.519 0.483 0.469 0.623 0.605
B.5 Prometheus-2-7B Decode 0.845 0.847 0.512 0.493 0.488 0.480 0.519 0.483 0.591 0.576
TRACT (ours)
6 C-RAFT Self C-RAIL 0.931 0.930 0.518 0.501 0.593 0.552 0.555 0.529 0.650 0.628
Ablation analysis for TRACT
A.1 C-RAFT GPT-4 C-RAIL 0.879 0.880 0.418 0.419 0.528 0.513 0.399 0.418 0.556 0.558
A.2 CE Self C-RAIL 0.919 0.917 0.468 0.436 0.562 0.526 0.517 0.503 0.617 0.596
A.3 CE Self Decode 0.873 0.873 0.358 0.346 0.418 0.404 0.435 0.426 0.521 0.512
A.4 TRACT with stage 2 initialized from $p_{\rm s}$ 0.674 0.684 0.448 0.437 0.505 0.477 0.432 0.421 0.515 0.505
Table 2: The results for Mistral-7B-Instruct. The best and second-best results (excluding ablations) for each column are marked with boldface and underline, respectively. Explanation of abbreviations: Train: training objective; Data: source of CoT used for training; Inf.: inference method; FB Bench: Feedback Bench; Vic. Bench: Vicuna Bench; $r$: Pearson’s correlation coefficient; $\rho$: Spearman’s rank correlation coefficient; C-RAFT: CoT-RAFT; C-RAIL: CoT-RAIL. $\dagger$: Prometheus-2-7B is obtained by merging two models trained from Feedback Collection and Preference Collection. We re-run the inference using the model released by Kim et al. (2024b) with the official code. $\ddagger$: When training without CoTs, the training target is simply the score, which is still generated by GPT-4.
Training and Evaluation
We fine-tune the LLMs with LoRA (Hu et al., 2022) using Llama-Factory (Zheng et al., 2024). We set $\lambda=1$ in Equation 4 unless otherwise stated.
During inference, for each sample in a dataset, we predict a score from the LLM using either standard decoding or the predictor in Equation 1 or 3, and calculate the correlation coefficient against the ground truth scores in the datasets. We report Pearson’s $r$ (Pearson, 1895), Spearman’s $\rho$ (Spearman, 1961), and Kendall’s $\tau$ (Kendall, 1938) for evaluation (see Kendall’s $\tau$ results in Appendix B.2; we find the trends to be consistent with other metrics). We report all hyperparameters in Appendix A.
@@ 5 Main Results
TRACT outperforms standard training and decoding with CoT.
Standard training and decoding with CoT (row B.2) is the default training and inference for LLM-as-a-judge in most prior works (Kim et al., 2024a, b). For both Mistral and Llama, TRACT (row 6) outperforms this baseline by a significant margin. When the base model is Mistral, TRACT reaches Pearson’s $r$ as high as 0.650, significantly outperforming training and inference with CoT (row B.2), which only attains Pearson’s $r$ of 0.557. When using Llama, TRACT improves the average Pearson’s $r$ by 0.100.
TRACT outperforms prior equally-sized SoTA.
Across all datasets, TRACT, trained without additional data, consistently outperforms Prometheus-2-7B (row B.5), the prior SoTA of the same size, by an average of 0.059 Pearson’s $r$. While we use self-generated data, this differs from Prometheus-2-7B, which is further trained on Preference Collection, a dataset generated by GPT-4.
TRACT outperforms RAFT.
RAFT (row B.4) can be considered as an ablation of TRACT that removes CoT. TRACT outperforms RAFT, showing that CoT helps the model predict the score. Quite surprisingly, RAFT achieves the best performance among all baselines. RAFT even outperforms Prometheus-2-7B (row B.5) by 0.032 Pearson’s $r$, which is trained over additional data corresponding to a ranking task. The same observation holds for the models trained using Llama-3.1.
@@
Training on self-generated CoTs is critical.
Row A.1 corresponds to training with CoT-RAFT on the GPT-4 generated data, and running inference with CoT-RAIL, without further training on self-generated CoTs. The average Pearson’s $r$ without training on self-generated data is 0.094 lower than TRACT. This shows that self-generated CoT is an important component of TRACT.
Using CE for training harms the performance.
Row A.2 is the result when we replace the fine-tuning objective from CoT-RAFT with the standard CE loss but still use CoT-RAIL in inference. This configuration leads to a Pearson’s $r$ 0.033 lower than TRACT. This shows that using CoT-RAFT for fine-tuning better aligns the fine-tuning objective and the predictor used during inference.
CoT-RAFT objective is necessary: self-generated CoTs alone are insufficient.
Prior research has demonstrated that fine-tuning LLMs on self-generated data using the CE loss can be beneficial (Yang et al., 2024). Here, we investigate whether CE loss fine-tuning on self-generated data alone can improve the performance compared with training on GPT-4 CoTs. Contrary to expectations, standard CE loss fine-tuning on self-generated CoTs (row A.3) yielded a significantly lower average Pearson’s $r$ of 0.512, far below TRACT’s performance. Moreover, this approach performed worse than fine-tuning on GPT-4 generated data (row B.2). This shows that self-generated CoTs are harmful when fine-tuned with CE loss but are beneficial when combined with the CoT-RAFT objective. Our finding marks the difference between our work and prior work using self-generated CoTs with CE loss, highlighting the distinct role of self-generated CoTs in TRACT.
Stage 2 training needs to be initialized from the seed LLM.
In Algorithm 1, the model in stage 2 is initialized from the seed LLM $p_{0}$. As an alternative, we explore initializing from $p_{\rm s}$, the model trained in stage 1, while keeping everything else the same as TRACT. The result shown in row A.4 achieves only 0.515 Pearson’s $r$. This shows that initialization in stage 2 from the seed LLM (as Algorithm 1 indicates) is crucial.

### v2

@@ Abstract
The LLM-as-a-judge paradigm uses large language models (LLMs) for automated text evaluation, which assigns a score to the text based on some scoring rubrics. Existing methods for LLM-as-a-judge use cross-entropy (CE) loss for fine-tuning, which neglects the numeric nature of score prediction. Recent work addresses numerical prediction limitations of LLM fine-tuning through regression-aware fine-tuning, which, however, does not consider chain-of-thought (CoT) reasoning for score prediction. In this paper, we introduce TRACT (Two-stage Regression-Aware fine-tuning with CoT), a method combining CoT reasoning with regression-aware training. The training objective of TRACT combines the CE loss for learning the CoT reasoning and the regression-aware loss for the score prediction. TRACT consists of two stages: first, a seed LLM is fine-tuned to generate CoTs; next, we retrain the seed LLM using the CoTs generated by the LLM trained in stage 1. Experiments across four LLM-as-a-judge datasets and two LLMs show that TRACT significantly outperforms existing methods. Extensive ablation studies validate the importance of each component in TRACT. [^1] All code and models can be found at https://github.com/d223302/TRACT.
@@
Cheng-Han Chiang NTU GICE Taiwan dcml0714@gmail.com                        Hung-yi Lee NTU GICE Taiwan hungyilee@ntu.edu.tw                        Michal Lukasik Google Research New York mlukasik@google.com
@@ 1 Introduction
Large Language Models (LLMs) have been applied to evaluate written based on the fine-grained evaluation rubrics specified in the input by outputting a score indicating the quality (Chiang and Lee, 2023a). The fine-grained evaluation rubrics define the criteria for scoring answers, and each instance to be evaluated can have a set of customized fine-grained evaluation rubrics. This LLM-as-a-judge paradigm has become the standard way to evaluate LLMs (Zheng et al., 2023; Dubois et al., 2024; Abdin et al., 2024; Lambert et al., 2025), and has applications across diverse scenarios (Yuan et al., 2024; Chiang et al., 2024). To induce the fine-grained assessment ability of LLMs, LLM-as-a-judge is typically trained with cross-entropy (CE) loss to predict chain-of-thought (CoT) reasoning about the evaluation, followed by a score (Kim et al., 2024a, b; Li et al., 2024).
@@
Although fine-tuning LLM-as-a-judge with CE loss is a well-established practice (Kim et al., 2024a, b), it has intuitive limitations when applied to predicting numerical targets. For example, given a ground truth score ’1’, placing high probability on the token ’5’ is penalized the same as placing high probability on the token ’2’, despite the fact that they induce very different errors in terms of numerical metrics such as squared error. To mitigate this, Lukasik et al. (2025) proposed regression-aware fine-tuning (RAFT; see §2.3), a method using squared error loss in fine-tuning LLMs.
RAFT has shown promising results on regression tasks; however, it does not incorporate CoT reasoning Wei et al. (2022); Kojima et al. (2022). Chiang and Lee (2023b) show that not using CoT can harm the performance of LLM-as-a-judge. This makes us ask: Can CoT reasoning be effectively integrated into regression-aware fine-tuning?
In this paper, we propose Two-stage Regression-Aware fine-tuning with CoT reasoning (TRACT), a method that combines the benefits of CoT reasoning with RAFT for improved numerical prediction abilities.. In stage 1, TRACT fine-tunes the seed LLM over both the CoT supervision and the ground truth scores in the original training data. TRACT’s fine-tuning objective (which we refer to as CoT-RAFT) is the sum of CE loss over the CoT, and the RAFT loss over score prediction (see Figure 1(a)). In stage 2, TRACT fine-tunes the seed LLM from the CoTs generated by the model trained in stage 1 using the same CoT-RAFT objective. This two-stage pipeline is for reducing the distribution between the seed LLM and the CoT supervision used for fine-tuning in stage 2 (see Figure 1(b)). We summarize TRACT in Algorithm 1 and contrast its formulation against baselines in Table 1.
We conduct experiments on five datasets and two LLMs and show that TRACT consistently outperforms baselines by a large margin, including Prometheus-2-7B (Kim et al., 2024b), which is the best model of equal size from prior works. Our contributions are as follows:
@@
    We show that TRACT, while trained on point-wise LLM-as-a-judge datasets, can work well on RewardBench (Lambert et al., 2024), a pairwise comparison dataset (§7).
  4. (iv)
@@ 3.2 Regression-Aware Fine-Tuning with CoT
CoT reasoning elicits a reasoning $s$ before outputting the score $y$. We propose and define the CR (CoT-RAIL) predictor as first sampling a CoT $\hat{s}$ conditioning on the input $x$, and then applying the RAIL predictor when conditioning on $x$ and $\hat{s}$,
@@ 3.3 Self-Generated Chain-of-Thoughts (CoTs)
When training with CoT-RAFT (Eq. 4), the score predictor is conditioned on the CoT generated by a target model, denoted as $\hat{s}\sim p_{\rm t}(\cdot\,|\,x)$. However, at inference time, CoT-RAIL (Eq. 3) relies on self-generated CoTs from the fine-tuned model $p$, i.e., $\hat{s}\sim p(\cdot\,|\,x)$. In our case, the CoTs used for training are from GPT-4 ($p_{\rm a}$) and may be very different from the CoTs generated by the fine-tuned model $p$. This creates a mismatch between the CoTs used during training and inference, which can harm the performance, as empirically shown in §5.
@@ 4 Experiment Setup
Models We fine-tune TRACT from two LLMs, Mistral-7B-Instruct-v0.2 (Jiang et al., 2023) and Llama-3.1-8B-Instruct (Grattafiori et al., 2024).
Training Dataset We follow Kim et al. (2024b) and use Feedback Collection (Kim et al., 2024a) as the training set. The training set contains roughly 100K samples. The whole dataset, including the samples to be evaluated, the evaluation responses, and ground truth scores, are generated by GPT-4. We follow the procedure in §3.3 to construct $D_{\rm self}$. We analyze the quality of the self-generated CoTs in Appendix B.1.
Test Datasets We use four datasets for point-wise LLM-as-a-judge, following Kim et al. (2024b).
@@
Baselines We compare TRACT against multiple baselines that differ in their training objectives and in whether they use CoT reasoning, including: (1) Standard fine-tuning and decoding without CoT2.2), (2) Standard fine-tuning and decoding with CoT3.1), (3) Zero-shot RAIL: Using RAIL predictor (Equation 1) on the seed LLM $p_{0}$, (4) RAFT2.3). A detailed comparison of the baselines is shown in Table 1.
@@
Train/Inference Configuration FB Bench FLASK Vic. Bench MT Bench Average Id CoT Train Data Inf. $r$ $\rho$ $r$ $\rho$ $r$ $\rho$ $r$ $\rho$ $r$ $\rho$ Baselines B.1 CE GPT-4 Decode 0.890 0.891 0.355 0.361 0.429 0.414 0.279 0.268 0.488 0.483 B.2 CE GPT-4 Decode 0.872 0.872 0.413 0.407 0.463 0.456 0.480 0.482 0.557 0.554 B.3 RAIL 0.197 0.175 0.200 0.149 0.281 0.165 0.309 0.216 0.247 0.176 B.4 RAFT GPT-4 RAIL 0.932 0.930 0.509 0.502 0.567 0.519 0.483 0.469 0.623 0.605 Not directly comparable models (Different Training Data) B.5 Prometheus-2-7B Decode 0.845 0.847 0.512 0.493 0.488 0.480 0.519 0.483 0.591 0.576 6 CLoud (Reward model) Decode 0.381 0.376 0.228 0.168 0.229 0.311 0.511 0.506 0.337 0.340 TRACT (ours) 7 C-RAFT Self C-RAIL 0.931 0.930 0.518 0.501 0.593 0.552 0.555 0.529 0.650 0.628 Ablation analysis for TRACT A.1 C-RAFT GPT-4 C-RAIL 0.879 0.880 0.418 0.419 0.528 0.513 0.399 0.418 0.556 0.558 A.2 CE Self C-RAIL 0.919 0.917 0.468 0.436 0.562 0.526 0.517 0.503 0.617 0.596 A.3 CE Self Decode 0.873 0.873 0.358 0.346 0.418 0.404 0.435 0.426 0.521 0.512 A.4 TRACT with stage 2 initialized from $p_{\rm s}$ 0.674 0.684 0.448 0.437 0.505 0.477 0.432 0.421 0.515 0.505
Table 2: The results for Mistral-7B-Instruct. The best and second-best results (excluding ablations) for each column are marked with boldface and underline, respectively. Explanation of abbreviations: Train: training objective; Data: source of CoT used for training; Inf.: inference method; FB Bench: Feedback Bench; Vic. Bench: Vicuna Bench; $r$: Pearson’s correlation coefficient; $\rho$: Spearman’s rank correlation coefficient; C-RAFT: CoT-RAFT; C-RAIL: CoT-RAIL. $\dagger$: Prometheus-2-7B is obtained by merging two models trained from Feedback Collection and Preference Collection. We re-run the inference using the model released by Kim et al. (2024b) with the official code. $\ddagger$: When training without CoTs, the training target is simply the score, which is still generated by GPT-4. $\ast$: CLoud (trained from Llama-3-8B) is a reward model; it is not designed for LLM-as-a-judge benchmarks (Section 7).
Training and Evaluation We fine-tune the LLMs with LoRA (Hu et al., 2022) using Llama-Factory (Zheng et al., 2024). We set $\lambda=1$ in Equation 4 unless otherwise stated.
During inference, for each sample in a dataset, we predict a score from the LLM using either standard decoding or the predictor in Equation 1 or 3, and calculate the correlation coefficient against the ground truth scores in the datasets. We report Pearson’s $r$ (Pearson, 1895), Spearman’s $\rho$ (Spearman, 1961), and Kendall’s $\tau$ (Kendall, 1938) (see Kendall’s $\tau$ results in Appendix B.2; we find the trends to be consistent with other metrics).
@@ 5 Main Results
TRACT outperforms standard training and decoding with CoT. Standard training and decoding with CoT (row B.2) is the default training and inference for LLM-as-a-judge in most prior works (Kim et al., 2024a, b). For both Mistral and Llama, TRACT (row 7) outperforms this baseline by a significant margin. When the base model is Mistral, TRACT reaches Pearson’s $r$ as high as 0.650, significantly outperforming training and inference with CoT (row B.2), which only attains Pearson’s $r$ of 0.557. When using Llama, TRACT improves the average Pearson’s $r$ by 0.100.
TRACT outperforms prior equally-sized SoTA. Across all datasets, TRACT, trained without additional data, consistently outperforms Prometheus-2-7B (row B.5), the prior SoTA of the same size, by an average of 0.059 Pearson’s $r$. While we use self-generated data, this differs from Prometheus-2-7B, which is further trained on Preference Collection, a dataset generated by GPT-4.
TRACT outperforms RAFT. RAFT (row B.4) can be considered as an ablation of TRACT that removes CoT. TRACT outperforms RAFT, showing that CoT helps the model predict the score. Quite surprisingly, RAFT achieves the best performance among all baselines. RAFT even outperforms Prometheus-2-7B (row B.5) by 0.032 Pearson’s $r$, which is trained over additional data corresponding to a ranking task. The same observation holds for the models trained using Llama-3.1.
@@
Training on self-generated CoTs is critical. Row A.1 corresponds to training with CoT-RAFT on the GPT-4 generated data, and running inference with CoT-RAIL, without further training on self-generated CoTs. The average Pearson’s $r$ without training on self-generated data is 0.094 lower than TRACT. This shows that self-generated CoT is an important component of TRACT.
Using CE for training harms the performance. Row A.2 is the result when we replace the fine-tuning objective from CoT-RAFT with the standard CE loss but still use CoT-RAIL in inference. This configuration leads to a Pearson’s $r$ 0.033 lower than TRACT. This shows that using CoT-RAFT for fine-tuning better aligns the fine-tuning objective and the predictor used during inference.
CoT-RAFT objective is necessary: self-generated CoTs alone are insufficient. Prior research has demonstrated that fine-tuning LLMs on self-generated data using the CE loss can be beneficial (Yang et al., 2024). Here, we investigate whether CE loss fine-tuning on self-generated data alone can improve the performance compared with training on GPT-4 CoTs. Contrary to expectations, standard CE loss fine-tuning on self-generated CoTs (row A.3) yielded a significantly lower average Pearson’s $r$ of 0.512, far below TRACT’s performance. Moreover, this approach performed worse than fine-tuning on GPT-4 generated data (row B.2). This shows that self-generated CoTs are harmful when fine-tuned with CE loss but are beneficial when combined with the CoT-RAFT objective. Our finding marks the difference between our work and prior work using self-generated CoTs with CE loss, highlighting the distinct role of self-generated CoTs in TRACT.
Stage 2 training needs to be initialized from the seed LLM. In Algorithm 1, the model in stage 2 is initialized from the seed LLM $p_{0}$. As an alternative, we explore initializing from $p_{\rm s}$, the model trained in stage 1, while keeping everything else the same as TRACT. The result shown in row A.4 achieves only 0.515 Pearson’s $r$. This shows that initialization in stage 2 from the seed LLM (as Algorithm 1 indicates) is crucial.